When 3.50 × 1022 molecules of ammonia react with 6.50 × 1022 molecules of oxygen according to the chemical equation shown below, how many grams of nitrogen gas are produced?

4 NH3(g) + 3 O2(g)→ 2 N2(g) + 6 H2O(g)



1
Expert's answer
2022-02-16T17:39:03-0500

Number of moles= Given molecules÷ Avogadro's number

Moles of Ammonia= (3.5×10^22)÷(6.023×10^23)=0.0581

Moles of Oxygen= (6.5×10^22)÷(6.023×10^23)=0.108

Ammonia is the limiting reagent as it limits formation of product:

4moles of Ammonia produces 2moles of Nitrogen.

0.0581 moles of Ammonia will produce=2/4 × 0.0581 = 0.0291 moles of nitrogen.

Molar mass of nitrogen= 28g/mol

Amount of nitrogen produced=

0.0291 × 28 = 0.813g

= 0.813g of nitrogen.